OpenStudy (anonymous):

Use an angle sum identity to verify the identity cos2θ = 2cos^2 θ-1.

OpenStudy (anonymous):

\[\cos 2\theta=\cos (\theta+\theta)=\cos \theta.\cos \theta-\sin \theta.\sin \theta=\cos ^{2}\theta-\sin ^{2}\theta=\cos ^{2}\theta-(1-\cos ^{2}\theta)=\cos ^{2}\theta-1+\cos^{2}\theta=2\cos^{2}-1\]

OpenStudy (anonymous):

and then comes your RHS